OpenStudy (anonymous):

At what rate (with respect to time) is the angle θ between the ground and the ladder changing, if the top of the ladder is sliding down the wall at the rate of r inches per second, at the moment that the top of the ladder is h feet from the ground? (You're looking here for an equation in terms of h and θ.)

OpenStudy (anonymous):

I know that dtheta/dt=dtheta/dh * dh/dt

OpenStudy (anonymous):

and that arcsin(h/25)=theta

OpenStudy (anonymous):

and the derivative of arcsin is equal to 1/sqrt(1-x^2)

OpenStudy (anonymous):

Pleas help...I just don't know how to put it altogether.

OpenStudy (nincompoop):

is this a general question without dimensions?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

can you draw what the problem is asking?

OpenStudy (anonymous):

I got : r/sqrt(1/(h/25)^2) but the answer choices say Otherwise :( I have the answer choices and picture attached

OpenStudy (nincompoop):



OpenStudy (nincompoop):

you have dimensions!

OpenStudy (anonymous):

yes, just the height of the ladder aka hypotenuse of the triangle

OpenStudy (anonymous):

...any ideas about the next step?

OpenStudy (nincompoop):

yeah

OpenStudy (anonymous):

I got r/sqrt(1/(h/25)^2) but I know I tripped cause its not in the answer choice. can you tell me where I missed up.

OpenStudy (nincompoop):

how did you get that is what I want to know

OpenStudy (anonymous):

dtheta/dt=dtheta/dh * dh/dt

OpenStudy (anonymous):

and that arcsin(h/25)=theta

OpenStudy (nincompoop):

:(

OpenStudy (anonymous):

I also know by the formula of the derivative of aarcsin that it is equal to 1/sqrt(1-x^2) and so I plugged in h/25 for x and multiplied everything by r

OpenStudy (nincompoop):

can you set it up using tangent?

OpenStudy (nincompoop):

tanθ = h/√(625-h^2)

OpenStudy (nincompoop):

secθ = 25/√(625-h^2)

OpenStudy (nincompoop):

do you know where to take from here?

OpenStudy (anonymous):

no, I got confuse. WHy is it in terms of secant instead of sine?

OpenStudy (nincompoop):

why sine?

OpenStudy (anonymous):

cause sintheta= (h/25)-->arcsin(h/25)=theta

OpenStudy (nincompoop):

how far does it get you?

OpenStudy (anonymous):

then I took the derivative of arcsin(h/25)=theta to get 1/sqrt(1/(h/25)^2)

OpenStudy (anonymous):

I multiplied that with r to get r/sqrt(1/(h/25)^2)

OpenStudy (anonymous):

and this was my final answer, but it is not on the answers. I attached them if you want to check them out.

OpenStudy (nincompoop):

you need to justify why you're doing every step

OpenStudy (nincompoop):

look at this tanθ = h/√(625-h^2)

OpenStudy (anonymous):

Your final answer, was it on the answer choices I gave?

OpenStudy (nincompoop):

it is not finished

OpenStudy (nincompoop):

dθ/dt = 1/√(625-h^2) dh/dt

OpenStudy (anonymous):

what do I do after that? can you give me the next step?

OpenStudy (anonymous):

A solution using Mathematica 9 is attached.